OpenStudy (anonymous):

What is the vertex of the parabola with the equation (y – 1)2 = 4x?

OpenStudy (amistre64):

since its not at the origin, we gotta drag it to the origin to look at it and examine it; how do we make: (x,y-1) = (0,0) ??

OpenStudy (anonymous):

-1

OpenStudy (amistre64):

x = 0 when x = ?? y-1 = 0 when y = ??

OpenStudy (amistre64):

x = 0 when x = 0 right? So the vertex so far as at (0,y) y-1 = 0 when y = ?? ; then we can see that the vertex is at (0,y)

OpenStudy (anonymous):

i think you meant (0,1)

OpenStudy (amistre64):

i did :)